Lisa wanted to find the actual distance from the top of the slide to the pool. The height of the slide is 8 feet and the distance from the bottom of the ladder to the pool is 6 feet.

Answer by rothauserc(4718)

we need to solve a right triangle for its hypotenuse(h)
:
8^2 + 6^2 = h^2
:
64 + 36 = h^2
:
h = square root(100)
:
h = 10
:
The actual distance from the top of the slide to the pool is 10 feet
: